OpenStudy (nicole143):

Could someone please explain to me how to solve this equation? (x-2)^3/4 = 8

OpenStudy (nicole143):

@nincompoop Could you help?

OpenStudy (anonymous):

\[(x-2)^{\frac{ 3 }{ 4 }}=8\] Multiply both sides with 4/3 exponent: \[((x-2)^{\frac{ 3 }{ 4 }})^{\frac{ 4 }{ 3 }}=(8)^{\frac{ 4 }{ 3 }}\] \[(x-2)=(2)^{4}\] \[x-2=16\] x=16+2 x=18

OpenStudy (nicole143):

Thank you soo much! I hadn't thought about multiplying by the reciprocal.

OpenStudy (anonymous):

you're welcome

OpenStudy (nicole143):

Could you help me with one more?

OpenStudy (nicole143):

For the function f(x) = (3 - 4x)^2, find f^-1. Also, find whether f^-1 is a function. I don't know how to deal with f^-1

OpenStudy (nicole143):

@jorea143

OpenStudy (anonymous):

That's inverse function. \[f(x) = (3 - 4x)^{2}\] \[y = (3 - 4x)^{2}\] Switch x and y: \[x = (3 - 4y)^{2}\] then solve for y: \[\pm \sqrt{x} = \sqrt{(3 - 4y)^{2}}\] \[\pm \sqrt{x} = 3 - 4y\] \[4y=3 \pm \sqrt{x}\] \[y=\frac{3 \pm \sqrt{x}}{4}\] Change y to \[f^{-1}(x)\] so \[f^{-1}(x)=\frac{3 \pm \sqrt{x}}{4}\]

OpenStudy (nicole143):

Thank you! I hope you have a good night!

OpenStudy (anonymous):

You're welcome. you too